If k ≥ 1, the graphs of y = sinx and y = keex
intersect for
x ≥ 0. Find the smallest value of k for which the graphs are tan￾gent.
k =
What are the coordinates of the point of tangency?
x = , y = .
1
Expert's answer
2020-03-16T11:03:54-0400

For tangency,

"sin x=ke^{-x}" and "cos x= -ke^{-x}"

"sinx=-cos x=ke^{-x}"

"tanx=-1 for x= n\u03c0+3\u03c0\/4"

"1/\sqrt{2}=ke^{-3π/4} ; x = 3π/4 ; n=0"

"k=\frac {e^{3π/4}}{\sqrt{2}}"

Coordinates of the point of tangency "(\frac{3π}{4},\frac 1{\sqrt{2}})"